OpenStudy (anonymous):

Find the Domain and Range of the following sets of ordered pair: {(x,y) | y-2x=4} {(x,y) | y=1-x/x-2} {(x,y) | y=x^2-2} {(x,y) | y=√ 4-x} Any Ideas guys?

OpenStudy (lgbasallote):

those are separate questions right?

OpenStudy (anonymous):

Nope. They are on the same number.

OpenStudy (anonymous):

Uhhh wait. Yeah they are. Just on the same number.

OpenStudy (anonymous):

It's like 1. a) b) etc.

OpenStudy (lgbasallote):

ohhh

OpenStudy (lgbasallote):

are you familiar with domain restrictions?

OpenStudy (anonymous):

Uhmm nope. I think we did study that before. But it' been a year since then. But I do know that the Domain is the X and the Range is Y. It's just that the set are quite complicated with all the roots and division going on.

OpenStudy (anonymous):

Some of them are even Quadratic Functions right?

OpenStudy (anonymous):

Let me just search for that first. So what does that got to do with the Sets?

OpenStudy (lgbasallote):

domain restrictions happen ONLY when: 1) \(\frac{1}{f(x)}\) f(x) means an expression with x variables 2) \(\sqrt{f(x)}\) same meaning of f(x) now look at a) is it in the form of either one of these?

OpenStudy (anonymous):

Yes? So I have to rewrite it into slope form with is y=mx+b?

OpenStudy (lgbasallote):

look at a...is there any fraction?

OpenStudy (anonymous):

NOpe

OpenStudy (lgbasallote):

is there any square root symbol?

OpenStudy (anonymous):

Nope. So does that mean. That it doesn't have Domain Restriction?

OpenStudy (lgbasallote):

yes..that means the domain for a is all real numbers

OpenStudy (lgbasallote):

now...to find range you have to solve for x...what do you think you should do to isolate x?

OpenStudy (lgbasallote):

we're still talking about a) btw

OpenStudy (anonymous):

Uhhhm. Move Y to the right side and then Divide both side with 2?

OpenStudy (lgbasallote):

how will it look like? can you type it?

OpenStudy (anonymous):

x=4+y/2

OpenStudy (lgbasallote):

now..is there a fraction?

OpenStudy (anonymous):

Yes

OpenStudy (lgbasallote):

is there a y in the denominator?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

There's 2

OpenStudy (lgbasallote):

therefore first restriction is crossed out...because there has to be a variable in the denominator is there a square root symbol?

OpenStudy (anonymous):

Nope

OpenStudy (lgbasallote):

therefore the range for a is all real numbers

OpenStudy (lgbasallote):

now...you may ask what if there is a restriction...

OpenStudy (anonymous):

Yeah

OpenStudy (lgbasallote):

let's take b for example

OpenStudy (lgbasallote):

is there a fraction?

OpenStudy (anonymous):

Yes there is, and there's also a variable

OpenStudy (anonymous):

in the denominator

OpenStudy (lgbasallote):

therefore there is arestriction...

OpenStudy (anonymous):

Yes

OpenStudy (lgbasallote):

to find this restriction equate the denominator to 0 (note this is only done for the first restriction) what do you get when you equate the denominator to 0?

OpenStudy (anonymous):

Uhmmm ..wait

OpenStudy (lgbasallote):

sure

OpenStudy (anonymous):

Does Y become 0?

OpenStudy (anonymous):

LIke 0 = 1-x/x-2?

OpenStudy (lgbasallote):

dont mind anything else...just the denomiantor

OpenStudy (anonymous):

Ohhh!

OpenStudy (anonymous):

It become the opposite

OpenStudy (anonymous):

x=2, x=-1

OpenStudy (lgbasallote):

just the \(\mathbf{denominator}\) x - 2 = 0 there fore x = 2 that is the restriction the domain is all real numbers except 2

OpenStudy (lgbasallote):

got it?

OpenStudy (anonymous):

Yep same goes for the Range right?

OpenStudy (lgbasallote):

for range...you have to solve for x first before attempting to find restircitions

OpenStudy (anonymous):

*domain sorry

OpenStudy (anonymous):

What happens with the Numerator? I don't touch it?

OpenStudy (lgbasallote):

but that is the domain

OpenStudy (lgbasallote):

when finding for restrictions in restriction #1 you just look at the denominator

OpenStudy (anonymous):

So every time I have a equation such as that. I equate the denominator only to get the Domain?

OpenStudy (lgbasallote):

yep

OpenStudy (anonymous):

Ohhh . Okay now I understand. So how do I solve for the Range?

OpenStudy (anonymous):

Since I already have X do I just substitute?

OpenStudy (lgbasallote):

that's usually the tricky part..start by cross multiplying

OpenStudy (lgbasallote):

no you dont sub...2 is not a value of x...it's the opposite actually...it's the not value of x...x is true for all numberS EXCEPT 2

OpenStudy (anonymous):

Okay. So what Do I cross multiply?

OpenStudy (lgbasallote):

y(x-2) = x - 1 got that?

OpenStudy (anonymous):

Ohhhhhhh. Okay . So I equate them?

OpenStudy (lgbasallote):

nope...distribute y what do you get??

OpenStudy (anonymous):

Ohh wait. so Multiply both side with the denominator to transfer it to the other side of the equation?

OpenStudy (lgbasallote):

note this kind of manipulation..it is very common ;)

OpenStudy (lgbasallote):

yep

OpenStudy (lgbasallote):

that's what we call cross multiply

OpenStudy (anonymous):

Okay. So after cross multiplying>

OpenStudy (lgbasallote):

distribute y into x - 2

OpenStudy (anonymous):

xy-2y=1-x

OpenStudy (anonymous):

Then?

OpenStudy (anonymous):

Solve for Y?

OpenStudy (anonymous):

1-x(-xy)/2 ?

OpenStudy (lgbasallote):

range means solving for x remembeR?

OpenStudy (anonymous):

Wait, Isn't Range Y?

OpenStudy (lgbasallote):

so xy - 2y = 1 - x put all x on one side..all non x to the other

OpenStudy (lgbasallote):

range is the set of y yes...look at when we were solving for domain..it was y = f(x) <---F(x) is an expression of x

OpenStudy (anonymous):

OHh okay.

OpenStudy (lgbasallote):

so put all x on one side?

OpenStudy (anonymous):

Yep it would look like. -1+y-2y= x-x ?

OpenStudy (anonymous):

or is it 2x = 1-y+2y?

OpenStudy (anonymous):

so Divide by 2 ir would be x= 1+y/2?

OpenStudy (lgbasallote):

it's actually neither

OpenStudy (anonymous):

what? Then what is it?

OpenStudy (lgbasallote):

im gonna save you the solving

OpenStudy (anonymous):

That would really help.

OpenStudy (lgbasallote):

put all the x on one side so it would just be xy +x = 1 - 2y

OpenStudy (lgbasallote):

got that?

OpenStudy (anonymous):

OHhhh. What I did i, that I removed X from XY! Silly me! So you can't do that right?

OpenStudy (anonymous):

Yep I got it

OpenStudy (anonymous):

But isn't 2y negative? So when you move it , it becomes positive right?

OpenStudy (lgbasallote):

now...you "remove" x by factoring it x(y +1) = 1 - 2y

OpenStudy (lgbasallote):

yeah my mistake

OpenStudy (lgbasallote):

x(y+1) = 1 + 2y

OpenStudy (anonymous):

Where did you get 1? x=1?

OpenStudy (anonymous):

Ohh now I get it

OpenStudy (lgbasallote):

you good?